From GlueXWiki
Jump to: navigation, search

November 3, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Clean room status
    • Production update
    • First package tests (Lubomir)
  2. Engineering update (Bill)
  3. Electronics update (?)
  4. Testing set-up at EEL126 (Beni)
  5. TDR (Lubomir)
  6. Other: naming convention


Participants: Bill, Dave, Vladimir, Beni, Mark, and Lubomir.


- Dave on production (see the link above for details): Six cathodes for package#2 ready. Scanning of wire positions: the hardware problem we have with the controller (ADC from the sensor gives numbers in 2V range instead of 10V) was "fixed" by reducing the threshold to 1.8V in the software. Surprisingly now we almost never miss wires; before we had ~10% of filed wires missed. Still, Dave will write a report about the problem that will be sent to Parker (with Larry's help) till the controller is still has warranty. Unfortunately, it may take several weeks to fix the problem and we don't want to do this now.

- Dave: yesterday we lost one phase at Blue Crab that affected the HVAC system for 2 hours. The particle counts went up to 41,000, but then dropped down to standard (~400) in few hours after the power came back. Bill suspects some dust accumulated in the returned ducks went into the clean room. Ron Bartek is aware of the problem. Fortunately only HVAC was affected and we could finish the electroplating that started before the outage, but we decided to install UPS system to be used for such critical operations. Mark offered a UPS and help to install it at Blue Crab. The UPS will be installed outside the clean room with extension cords going inside.

- Lubomir: After we found the noise on the top cell of the first package originating from two places on the wire frame (+HV traces on the back of the PCBs) we looked more carefully at the same places in the other cells. Similar noise was found on another two cells but with smaller amplitude and frequency. We found the frequency of the noise (and interestingly also the dark current) depends somehow on the pressure you apply on the PCB in this area by the torque on the nearby threated rod. Likely explanation is that we have some sparking between the HV traces and the rochacell due to possible scratches or de-lamination of the trace (Bill). In any case the frequency is no more than few hundred Hz which is not a problem.


- Bill showed a scheme for the positions of the signal connectors in the package; it will be linked to this/next meeting after some update. Related to that, once more we discussed the FDC naming convention linked above.

- Cooling system: need to build tool that will help forming the right tubing diameter (120cm), maybe out of plywood. Casey may do it tomorrow.

- Right after the meeting Bill installed the holding brackets on the first package.

- Gas system: following the discussing on the collaboration meeting yesterday, Lubomir made estimates about the oxygen diffusion into the gas volume due to the open end of the thin tubes used at the output of the cells. The diffusion length per second is less than a cm. The speed in the thin tube is at least 10cm/sec which at 2m length makes the oxygen concentration negligible. Today we tried using one bubbler after the thin tubes, it worked better than without thin tubes, but stopped bubbling after running for ~30 min at a flow of 80 ccpm. Needs further studies.


- Chris (not at the meeting) will work next week on the pre-amps: installing cooling brackets and testing; possibly Vladimir will help him.

Chamber testing at 126

- Beni: Will get two more VXS crates, need more 20Amp power sockets, Mark will help him.


- Lubomir started working on the FDC part of the TDR, current version will be at the link above, right now very pre-preliminary. Beni suggested to concentrate on the technical description of the detector and just refer to the CDR for the motivations and MC studies.